40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 6 18:14:33 2013, Manasa, Summary, LSC, Calibration of BS, ITMX and ITMY actuators MICHerrCalibration.pngMICHerrCalibration.pngMICHOLTF.pngactcalibITMBS.png
    Reply  Thu Mar 7 01:43:35 2013, yuta, Update, LSC, calibrated MI differential length spectra LSCMICHOLTF_MI.pngMImotion.png
    Reply  Fri Mar 8 02:17:04 2013, yuta, Update, LSC, calibration of PRM actuator calibREFL55Q.pngactcalibPRM.pngOLTF_PRCL.png
       Reply  Fri Mar 8 03:07:19 2013, yuta, Update, LSC, calibrated PRM-ITMY length spectra LSCPRCLOLTF_PRITMY.pngPRITMYmotion.png
Message ID: 8255     Entry time: Fri Mar 8 02:17:04 2013     In reply to: 8242     Reply to this: 8256
Author: yuta 
Type: Update 
Category: LSC 
Subject: calibration of PRM actuator 

[Manasa, Yuta]

We measured AC response of PRM actuator using PRM-ITMY cavity.
Result is

PRM:  (19.6 +/- 0.3) x 10^{-9} (Hz/f)^2 m/counts

It is almost the same as in 2011 (elog #5583). We took the same procedure as what Kiwamu did.

What we did:
  1. Aligned PRMI in usual procedure, mis-aligned ITMX and locked PRM-ITMY cavity using REFL55_Q_ERR. POP DC was about 18 when locked.

  2. Made UGF of PRM-ITMY cavity lock at 10 Hz and introduced elliptic LPF at 50 Hz(OLTF below).
OLTF_PRCL.png


  3. Measured transfer function from C1:LSC_ITMY_EXC to C1:LSC_REFL55_Q_ERR. Dividing this by ITMY actuator response(measured in elog #8242) gives calibration of REFL55_Q.

  4. Measured transfer function from C1:LSC_PRM_EXC to C1:LSC_REFL55_Q_ERR to calibrate PRM actuator.

Result:
  Calibration factor for REFL55_Q for PRM-ITMY cavity was (1.37 +/- 0.02) x 10^9 counts/m (plot below). Error is mainly from statistical error of the average.
calibREFL55Q.png


  Measured AC response (50-200 Hz) of PRM is below.
actcalibPRM.png


Next:
  - Measure free-run length spectrum of PRM-ITMY cavity and compare with MICH free-run.

ELOG V3.1.3-